OpenStudy (anonymous):

Tom invests $10,000 in a savings account that offers 3.5 percent interest, compounded continuously. In 10 years Tom will have earned $ in interest, and in years the investment will double. (Use the rule of 70 where required.)

OpenStudy (retireed):

Part 1) \[A = Pe ^{rt}\] P = 10,000 r = 0.035 t = 10 Interest = A - P

OpenStudy (retireed):

and in t years the investment will double. (Use the rule of 70 where required.) A = Pe^rt where A = 2P 2P = Pe^rt divide both sides by P ln(2) = rt t = ln(2) / r Solve for t when r = 0.035